The Bistro is planning to add a new line of noodles that will require the acquisition of new processing equipment. The equipment will cost $1,000,000, including installation and shipping. It will be depreciated straight-line to zero value over the 10-year economic life of the project. Interest cost associated with financing the equipment purchase is estimated to be $40,000 annually. The expected salvage value of the machine at the end of 10 years is $200,000. One year ago a marketing survey was performed to gauge the likely success of this new project. The survey cost $25,000, and was paid last year. If this new equipment is acquired, it will also allow the replacement of old equipment used for other food lines. This old equipment can be salvaged for $150,000 and has a book value of $200,000. The remaining depreciation on the old equipment is $40,000 annually for five more years. Additional net working capital of $85,000 will be needed immediately. When the project is terminated in 10 years, there no longer will be a need for this incremental working capital. The Bistro expects to sell $300,000 worth of this new pasta annually. The cost of producing and selling the pasta is estimated to be $50,000 annually (not including depreciation or interest expense). The marginal tax rate is 40 percent. 

Calculate all cash flows for the project:

Year 0 Cash flow=______

Year 1-5 cash flow=______

Year 6-9 cash flow=______

Year 10 cash flow=______
Answer #1

Given:

The old equipment can be salvaged for $150,000 and has a book value of $200,000.

Tax benefit on loss = (200000 - 150000) * 40% = $20,000

Salvage value including tax benefit = 150000 + 20000 = $170,000

The survey cost of $25,000 was paid last year. The same is a sunk cost and not relevant for this analysis.

Required year wise cash flows are calculated and given (highlighted yellow) below:

Year 1. 4 10 New Equipment cost Salvage value of old equipment with tax benefit Net working capital ($1,000,000) $170,000 ($8